SALES INCREASE Paladin Furnishings generated $4 million in sales during 2016, and its year-end total assets were $3.2 million. Also, at year-end 2016, current liabilities were $500,000, consisting of $200,000 of notes payable, $200,000 of accounts payable, and $100,000 of accrued liabilities. Looking ahead to 2017, the company estimates that its assets must increase by $0.80 for every $1.00 increase in sales. Paladin’s profit margin is 3%, and its retention ratio is 50%. How large of a sales increase can the company achieve without having to raise funds externally?
Answer #1

AFN = (A/S0)ΔS–(L/S0)ΔS–MS1(RR)

A- Assets tied directly to sales

L-spontaneous liabilities that are affected by sales

S0=the previous year's sales

S1=total projected sales for next year

ΔS=the change in sales between S0 and S1

MS1=projected net income

RR=the retention ratio from net income

$0 = ($3,200,000/$4,000,000)×(S1-S0)–(($200,000+$100,000)/$4,000,000)×(S1-S0)–S1×3%×50%

$0 = ($3,200,000/$4,000,000)×(S1-$4,000,000)–($300,000/$4,000,000)×(S1-$4,000,000)–S1×1.50%

$0 = 0.80×S1-$3,200,000–0.075×S1+$300,000–S1×1.50%

$2,900,000 = S1×(0.80-0.075-0.015)

S1 = $4,084,507

Maximum sales increase = $84,507 ($4,084,507-$4,000,000)